# Accounting for multiplicity in machine learning benchmark performance

### v4

@@
1 Introduction: Population vs sample state-of-the-art performance
@@
Machine learning methods are commonly evaluated and compared by their performance on data sets from public repositories, where methods can be evaluated under identical conditions and across time. Theoretical results play an important role in machine learning research, but for the last couple of decades the success of increasingly complex methods has pushed the proof towards the pudding: performance on benchmark data sets.
@@
The numbers don’t lie: if method A performs better than method B on a specific data set, that is what it did. But the truth that the numbers tell is limited. We think of a data set, for example a collection of images, as a random sample from an underlying population. In science we are interested in the performance on the population, and the observed performance acts as an estimate. A good estimator gives a reliable result, but overfitting to the data set, originating from wrong use of cross-validation, data leakage and data reuse, can give overly optimistic estimates. An independent test set, often referred to as a hold-out set, is a sample that is not used for development of the method, and gives an unbiased estimator. Competitions and challenges with prize money typically withhold the test set so that it becomes impossible for the participants to use it for development.
Although we know these things we have a tendency to extrapolate the performance metric and interpret it as something about the world rather than something about a data set. If we always appended “on MNIST” when we said so-and-so method had accuracy so-and-so nobody would care. Who cares about MNIST? We hold out data precisely because we want to estimate some sort of long-term behavior on similar data.
An unbiased estimator solves only half of the problem; the variance tells just as much as the bias regarding the quality of a method. For machine learning methods, the sources of variance are plentiful. One is the random nature of the test set, the training set and the validation set. In addition comes the setting of hyperparameters, which can lead to substantial variation, and minor variance sources such as random seeds and stochastic dropouts. The sources of variance that are part of the training are often referred to as influencing the method’s stability or robustness. With knowledge about the behaviour of both the bias and the variance of an estimator, the error can be predicted, often in terms of a confidence interval. It is certainly more useful to know what the lower limit of the 95% confidence interval is, than the observed performance on a specific test set.
@@
The highest ranked performance on a problem is referred to as state-of-the-art (SOTA) performance, and is used, among other things, as a reference point for publication of new methods. As a specific method’s performance on a test set is just an estimate of its true underlying performance, the reported SOTA performances are estimates, and hence come with bias and variance.
State-of-the-art performance increasingly relies on benchmark data sets, often in combination with a challenge or a competition. The practice of hold-out test sets in competitions safe-guards against overoptimistic estimates for the individual classifiers. Data is often released after the competition has ended, and new methods can be compared to the state-of-the-art performance.
@@
As we will show in the subsequent sections, using the highest-ranked performance as an estimate for state-of-the-art is a biased estimator, giving overoptimistic results. The mechanisms at play are those of multiplicity, a topic that is well-studied in the context of multiple comparisons and multiple testing, but has, as far as the authors are aware of, been nearly absent from the discussion regarding state-of-the-art estimates.
@@
Biased state-of-the-art has consequences both for practical implementations and for research itself. A health care provider might want to use some machine learning method for tumour detection, but only if the method correctly classifies at least 80%. The health care provider closely follows updates on state-of-the-art performance, and are ready to invest time and money in a pilot project once the lower bound surpasses 80%. If the state-of-the-art performance is an optimistic estimate, the health care provider has now lost time and money, and perhaps also some faith in research results.
@@
For the research community, the situation is perhaps even more dramatic. The optimistic state-of-the-art estimate is used as a standard for evaluating new methods. Not to say that a new method necessarily needs to beat the current state-of-the-art to catch interest from reviewers and editors, but it certainly helps if its performance is comparable. It has been pointed out (see Section 2) that the incremental increase in performance can obstruct the spread of new ideas that have substantial potential, but just haven’t reached their peak yet.
@@
In this article, we provide a probability distribution for the case of multiple classifiers. We demonstrate the impact of multiplicity through simulated examples. We show how classifier dependency impacts the variance, but also that the impact is limited when the accuracy is high. Finally, we discuss a real-world example; a Kaggle competition from 2020. All code is available at https://github.com/3inar/ninety-nine.
@@ 2 Related work
Multiple comparisons is a well-studied subject within statistics, see, e.g., Tukey (1991) for an interesting discussion on some topics of multiple comparisons, among them the confidence intervals. For a direct demonstration of how multiple testing can lead to false discoveries, Bennett et al. (2009) gives an example where brain activity is detected in a (dead) salmon “looking” at photos of people. In the machine learning context, with multiple data sets and classifiers, Demšar (2006) suggests adjusting for multiple testing, according to well-known statistical theory. Our article focuses on the estimate itself and its uncertainty, and not comparison of methods.
@@
The re-use of benchmark datasets has been addressed in several publications. Salzberg (1997) point to the problem that arises because these datasets are static, and therefore run the risk of being exhausted by overfitting. Exhausted public data sets still serve important functions, e.g., a new idea can show to perform satisfactory, even when it is not superior. Thompson et al. (2020) distinguishing between simultaneous and sequential multiple correction procedures. In sequential testing, which is the reality of public datasets, the number of tests is constantly updated. An interesting topic they bring up is the conflict between sharing incentive, open access and stable false positive rate. There are several demonstrations of the vulnerability of re-use of benchmark datasets. Teney et al. (2020) performed a successful attack on the VQA-CP benchmark dataset, that is, they are able to attain high performance with a non-generalisable classifier. They point towards the risk of ML methods capturing idiosyncrasies of a dataset. Torralba and Efros (2011) introduced the game Name That Dataset! motivating their concern of methods being tailored to a dataset rather than a problem. The lack of statistically significant differences among top performing methods is demonstrated by, e.g., Everingham et al. (2015) and Fernández-Delgado et al. (2014).
@@
The same mechanisms of overfitting are present in challenges through multiple submissions to the validation set. Blum and Hardt (2015) introduced an algorithm that releases the public leaderboard score only if there has been significant improvement compared to the previous submission. This will prevent overfitting by hindering adjustments to random fluctuations in the score, and not true improvements of the method. Dwork et al. (2017) use the principles of differential privacy to prevent overfitting to the validation set.
@@
Several publications show that the overfitting is not as bad as expected, and provide possible explanations. Recht et al. (2019) constructed new, independent test sets of CIFAR-10 and ImageNet. They found that the accuracy drop is dramatic, around 10%. However, the order of the best performing methods is preserved, which can be explained by the fact that multiple classes (Feldman et al. (2019)) and model similarity (Mania et al. (2019)) slows down overfitting. Roelofs et al. (2019) compared the results from the private leaderboard (one submission per team) and the public leaderboard (several submissions per team) in Kaggle challenges. Several of their analyses indicate overfitting even though their main conclusion is that there is little evidence of overfitting.
@@
Overfitting and its consequences are avoided in most challenges by withholding the test set so that a method can only be evaluated once and not adapted to the test set performance. Hold-out datasets also prevents wrong use of cross-validation, see, e.g., Fernández-Delgado et al. (2014) and the seminal paper of Dietterich (1998). Another contribution from public challenges is the requirement of reproducibility to collect the prize money, which is a well-known concern in science, see, e.g., Gundersen and Kjensmo (2018). Public challenges typically rank the participants by a (single) performance measure, and can fall victims of Goodhart’s law, paraphrased as “When a measure becomes a target, it ceases to be a good measure.” by Teney et al. (2020). Ma et al. (2021) offers a framework and platform for evaluation of NLP methods, where other aspects such as memory use and robustness are evaluated.
@@
The robustness/stability of a method, and the variability in performance are crucial aspects that the hold-out test set of public challenges does not offer an immediate solution to. Bouthillier et al. (2021) showed empirically that hyperparameter choice and the random nature of the data are two large contributors to variance, whereas data augmentation, dropouts, and weights initialisations are minor contributors. Choice of hyperparameters are often adjusted to validation set results, and the strategies of Blum and Hardt (2015) and Dwork et al. (2017) can hinder that, if implemented by the challenge hosts. Our contribution is a focus on better estimate for state-of-the-art performance under the current conditions, where the approach of Bouthillier et al. (2021) does not apply. Bousquet and Elisseeff (2002) investigate how sampling randomness influences accuracy, both for regression algorithms and classification algorithms, relying on the work of Talagrand (1996) who stated that “A random variable that depends (in a “smooth” way) on the influence of many independent variables (but not too much of any of them) is essentially constant.”
@@
There is a rich literature on various aspects regarding the use of public challenges as standards for scientific publications, see, e.g., Varoquaux and Cheplygina (2022) for a much broader overview than the one given here. Concerns and solutions regarding multiplicity and better performance estimates are active research fields. We contribute to this conversation by shining the light on how multiplicity of classifiers’ performance on hold-out test sets create optimistic estimates of state-of-the-art performance.
@@ 3.1 Notation
$\hat{\theta}(X)=\frac{n-X}{n}$ $:\text{estimator for }\theta,\text{ unbiased}$
@@
$\hat{\theta}^{\,max}_{SOTA}(\mathbf{X})=\max_{j}\hat{\theta}_{j}$ $:\text{estimator for }\theta_{SOTA},\text{ biased}$
@@
3.2 Two coin-flip examples
@@
Multiplicity is, simply put, how the probability of an outcome changes when an experiment is performed multiple times. Consider $n=20$ flips of a fair coin where the outcome is the number of heads, referred to as the number of successes in a binomial distribution, $\mathcal{B}(n,\theta)$, where the probability of success is $\theta=0.5$. Let the random variable $X$ denote the number of failures, $X\sim\mathcal{B}(n,1\text{-}\theta)$. The probability of $15$ heads or more corresponds to observing $x=n-15$ tails or less, that is, the cumulative distribution $P(X\leq x|n,1\text{-}\theta)$. This corresponds to the cumulative distribution for the estimator $\hat{\theta}(X)=\frac{n\text{-}X}{n}$, and we have that $P(\hat{\theta}(X)\geq\frac{n\text{-}x}{n})=P(X\leq x)=0.02069$ for $\hat{\theta}\geq 0.75$. In other words, if $\theta=0.5$, there is a probability of $0.02069$ that the estimate is $0.75$ or above in a series of $n=20$ trials. The probability of exactly $x=5$ can be calculated from the probability mass function (pmf) $P(X=x|n,1\text{-}\theta)=0.01479$, but the interest usually lies with the cumulative distribution. We focus on the case where $X\leq x$, but it is straightforward to substitute $P(X\leq x)$ for $P(X=x)$ in the following.
If we fix $x$ we can define the experiment described above as a Bernoulli trial, where success is defined as $X\leq x$, with corresponding probability $P(X\leq x)$, which we from now on write as $P_{x}$ for ease of notation. If we perform the coin-flip experiment $m$ times, and success is defined as $X\leq x$, we have a series of $m$ Bernoulli trials, and the number of successes follow the binomial distribution $\mathcal{B}(m,P_{x})$. Let $C_{x}$ be the random variable that denotes the number of successful experiments. The probability of, for example, $C_{x}=10$ experiments with at most $x=5$ tails, is calculated from the pmf $P(C_{x}=10|m,P_{x})$. Of special interest is the case where at least one experiment succeeds, $P(C_{x}>0)=1-P(C_{x}=0)$. In our example, if the coin-flip experiment is performed $m=100$ times, the probability of observing $15$ heads or more at least once is $P(C_{x}>0|m,P_{x})=1-P(C_{x}=0|m,P_{x})$, with $P_{x}=0.02069$ from above. In the estimation context, we have that although the probability of $\hat{\theta}\geq 0.75$ is only $0.02069$ for the individual experiments, the probability of at least one $\hat{\theta}\geq 0.75$ is $0.8765$ when the experiment is performed $m=100$ times.
@@
Consider an experiment where a coin is flipped $n=20$ times, and the number of successes is the number of times a classifier correctly predicts the outcome. An unbiased estimator for a classifier’s probability of correct prediction, $\theta$, is ${\hat{\theta}=(n-X)/n}$, where $X$ is the number of failures. The corresponding estimate is referred to as accuracy. The true state-of-the-art for coin-flip classifiers is $\theta_{SOTA}=0.5$, since it is impossible to train a classifier to predict coin flips. If $m=1,000$ classifiers independently give their predictions for $n=20$ flips, we can consider this as a binomial experiment with $1,000$ Bernoulli trials, where the probability of success is the individual classifier’s probability of $n-X\leq n-x$ correct predictions. The probability of at least one classifier achieving an accuracy of at least $0.90$ is $0.18231-P(C_{x}=0|m,P_{x})=$ 0.1823, where $P_{x}=P(X\leq 18|n,1\text{-}\theta)=0.00020$. This means that using the top-ranked method as an estimate and ignoring multiplicity there is a substantial risk of overestimating the state-of-the-art.
3.3 The probability distribution of $\hat{\theta}^{\,max}_{SOTA}$
We define the SOTA probability of correct prediction, $\theta_{SOTA}$, as the highest probability of correct prediction among all classifiers applied to a test set, in accordance with the common set-up for public competition. Of interest is $P(\hat{\theta}^{\,max}_{SOTA}|m,n,\theta)$, the probability of $\hat{\theta}^{\,max}_{SOTA}$, being the estimated state of the art from taking the highest ranked classifier, given the number of classifiers, the size of the test set, and the probabilities of correct prediction, $\theta=\{\theta_{1},\ldots,\theta_{m}\}$.
We will in the following assume independent classifiers, and discuss dependency in Section 4. We demonstrate the impact of multiplicity in the simple case of identical probabilities of correct prediction, $\theta_{j}=\theta$ for all classifiers, $j=1,\ldots,m$, and where $\theta_{SOTA}=\theta$. We discuss non-identical probabilities in Section 4.2.

### v5

@@
1 Estimating state-of-the-art performance
In this article, we direct attention to the reporting of state-of-the-art classifier performance by presenting the sample maximum estimator distribution function and analysing its bias. This analysis is done using the framework of multiplicity.
Within the fields of research that develop, refine, test and use classifiers, the term state-of-the-art (SOTA) commonly refers to the highest reported performance to date, and is used as a reference point when presenting and comparing models. The reported performance of a classifier, originating from applying that classifier to a specific data set, a sample, is in effect an estimate of that classifier’s underlying population performance.
The highest reported performance among a number of classifiers serves as an estimate of SOTA, although the term ’estimate’ is seldom used explicitly in the SOTA context. The highest reported performance corresponds to the sample maximum of population performances, and this estimator is biased when used to estimate SOTA. The sample maximum gives in overoptimistic estimates of the maximum population performance, and is problematic as a reference point in research. The mechanisms creating this bias comes from multiplicity, and the fact that the we pick the highest sample performance among multiple classifier results.
In its simplest form, with independent binary classifiers with identical probability of success, the distribution function of the sample maximum estimator exists in closed form with only a few parameters, and can be directly applied to calculate and analyse the bias in a certain situation. With growing complexity, such as mutual dependencies and more elaborate performance measures, the distribution function of the estimator is no longer known, but the mechanisms at play and the pitfalls to be aware of are much the same, and limited studies can be done by straightforward simulations.
@@ 1 Estimating state-of-the-art performance
Machine learning models are commonly evaluated and compared by their performance on data sets from public repositories, where models can be evaluated under identical conditions and across time. Theoretical results play an important role in machine learning research, but for the last couple of decades the success of increasingly complex models has pushed the proof towards the pudding: performance on benchmark data sets.
@@
The numbers don’t lie: if model A performs better than model B on a specific data set, that is what it did. But the truth that these numbers tell is limited. We are interested in the population performance, and the sample performance is only an estimate. A good estimating procedure gives a reliable result, and there are plenty of pitfalls to be aware of: Overfitting to the data set, originating from wrong use of cross-validation, data leakage and data reuse, can give overly optimistic estimates. An independent test set, often referred to as a hold-out set, is a sample that is not used for development of the model, and prevents overfitting. Competitions and challenges with prize money typically withhold the test set so that it becomes impossible for the participants to use it for training.
@@
The highest ranked performance on a problem is referred to as SOTA performance, and is used, among other things, as a reference point for publication of new models. As a specific model’s performance on a test set is just an estimate of its true underlying performance, the reported SOTA performances are estimates, and hence come with bias and variance. The practice of hold-out test sets in competitions safe-guards against overoptimistic estimates for the individual classifiers. Data is often released after the competition has ended, and new methods can be compared to the SOTA performance.
@@
As we will show in the subsequent sections, using the highest-ranked performance as an estimate for SOTA is a biased estimator, giving overoptimistic results. The mechanisms at play are those of multiplicity, a topic that is well-studied in the context of multiple comparisons and multiple testing, but has, as far as the authors are aware of, been nearly absent from the discussion regarding SOTA estimates.
@@
Biased SOTA estimation has consequences both for practical implementations and for research itself. A health care provider might want to use some machine learning method for tumour detection, but only if the method correctly classifies at least 80%. They follow updates on SOTA performance, and are ready to invest time and money in a pilot project once the lower bound surpasses 80%. If the SOTA performance is an optimistic estimate, the health care provider has now lost time and money, and perhaps also some faith in research results.
@@
For the research community, the situation is perhaps even more dramatic. The optimistic SOTA estimate is used as a standard for evaluating new methods. Not to say that a new method necessarily needs to beat the current SOTA to catch interest from reviewers and editors, but it certainly helps if its performance is comparable. It has been pointed out (see Section 2) that the incremental increase in performance can obstruct the spread of new ideas that have substantial potential, but just haven’t reached their peak yet.
@@
In this article, we provide a probability distribution for the case of multiple classifiers. We demonstrate the impact of multiplicity through simulated examples. We show how classifier dependency impacts the variance, but also that the impact is limited when the probability of success is high. Finally, we discuss three real-world examples that each highlights different aspects multiplicity and SOTA estimation. All code is available at https://github.com/3inar/ninety-nine.
@@ 2 Related work
Multiple comparisons is a well-studied subject within statistics, see, e.g., Tukey (1991) for an interesting discussion on some topics of multiple comparisons, among them the confidence intervals (CIs). For a direct demonstration of how multiple testing can lead to false discoveries, Bennett et al. (2009) give an example where brain activity is detected in (dead) salmon “looking” at photos of people. In the machine learning context, with multiple data sets and classifiers, Demšar (2006) suggests adjusting for multiple testing, according to well-known statistical theory. Our article focuses on the estimate itself and its uncertainty, and not comparison of methods.
@@
The re-use of benchmark datasets has been addressed in several publications. Salzberg (1997) point to the problem that arises because these datasets are static, and therefore run the risk of being exhausted by overfitting. Exhausted public data sets still serve important functions, e.g., a new idea can show to perform satisfactory, even when it is not superior. Thompson et al. (2020) distinguish between simultaneous and sequential multiple correction procedures. In sequential testing, which is the reality of public datasets, the number of tests is constantly updated. An interesting topic they bring up is the conflict between sharing incentive, open access and stable false positive rate. There are several demonstrations of the vulnerability of re-use of benchmark datasets. Teney et al. (2020) performed a successful attack on the VQA-CP benchmark dataset, that is, they are able to attain high performance with a non-generalisable classifier. They point towards the risk of ML methods capturing idiosyncrasies of a dataset. Torralba and Efros (2011) introduced the game Name That Dataset! motivating their concern of methods being tailored to a dataset rather than a problem. The lack of statistically significant differences among top performing methods is demonstrated by, e.g., Everingham et al. (2015) and Fernández-Delgado et al. (2014).
@@
The same mechanisms of overfitting are present in challenges through multiple submissions to the validation set. Blum and Hardt (2015) introduced an algorithm that releases the public leaderboard score only if there has been significant improvement compared to the previous submission. This will prevent overfitting by hindering adjustments to random fluctuations in the score and not true improvements of the method. Dwork et al. (2017) used the principles of differential privacy to prevent overfitting to the validation set.
@@
Several publications show that the overfitting is not as bad as expected, and provide possible explanations. Recht et al. (2019) constructed new, independent test sets of CIFAR-10 and ImageNet. They found that the accuracy drop is dramatic, around 10%. However, the order of the best performing methods is preserved, which can be explained by the fact that multiple classes (Feldman et al. (2019)) and model similarity (Mania et al. (2019)) slow down overfitting. Roelofs et al. (2019) compared the results from the private leaderboard (one submission per team) and the public leaderboard (several submissions per team) in Kaggle challenges. Several of their analyses indicate overfitting even though their main conclusion is that there is little evidence of overfitting.
@@
Overfitting and its consequences are avoided in most challenges by withholding the test set so that a method can only be evaluated once and not adapted to the test set performance. Hold-out datasets also prevents wrong use of cross-validation, see, e.g., Fernández-Delgado et al. (2014) and the seminal paper of Dietterich (1998). Another contribution from public challenges is the requirement of reproducibility to collect the prize money, which is a well-known concern in science, see, e.g., Gundersen and Kjensmo (2018). Public challenges typically rank the participants by a (single) performance measure, and can fall victims of Goodhart’s law, paraphrased as “When a measure becomes a target, it ceases to be a good measure.” by Teney et al. (2020). Ma et al. (2021) offered a framework and platform for evaluation of NLP methods, where other aspects such as memory use and robustness are evaluated.
@@
The robustness/stability of a method, and the variability in performance are crucial aspects that the hold-out test set of public challenges does not offer an immediate solution to. Bouthillier et al. (2021) showed empirically that hyperparameter choice and the random nature of the data are two large contributors to variance, whereas data augmentation, dropouts, and weights initialisations are minor contributors. Choice of hyperparameters are often adjusted to validation set results, and the strategies of Blum and Hardt (2015) and Dwork et al. (2017) can hinder that, if implemented by the challenge hosts. Our contribution is a focus on better estimate for SOTA performance under the current conditions, where the approach of Bouthillier et al. (2021) does not apply. Bousquet and Elisseeff (2002) investigated how sampling randomness influences accuracy, both for regression algorithms and classification algorithms, relying on the work of Talagrand (1996) who stated that “A random variable that depends (in a “smooth” way) on the influence of many independent variables (but not too much of any of them) is essentially constant.”
@@
There is a rich literature on various aspects regarding the use of public challenges as standards for scientific publications, see, e.g., Varoquaux and Cheplygina (2022) for a much broader overview than the one given here. Concerns and solutions regarding multiplicity and better performance estimates are active research fields. We contribute to this conversation by shining the light on how multiplicity of classifiers’ performance on hold-out test sets create optimistic estimates of SOTA performance.
@@ 3.1 Notation
$\hat{\theta}(X)=\frac{n-X}{n}$ $:\text{estimator for }\theta$
$\hat{\theta}=\frac{n-x}{n}$ $:\text{estimate of }\theta$
@@
$\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ $:\text{number of failures for classifiers }j=1,\ldots,m$
$C_{x}$ $:\text{number of classifiers with at most $x$ failures}$
$\,\,\,\,C_{x}\sim\mathcal{B}(m,P_{x})$
[…]
$\,\,\,\,z=0,1,2,\ldots,n$
@@
[…]
3.2 A coin-flip example
Multiplicity is, simply put, how the probability of an outcome changes when an experiment is performed multiple times.
@@ 3.2 A coin-flip example
Consider an experiment where a fair coin is flipped $n=20$ times and the number of successes is the number of times a classifier correctly predicts the outcome. Let the random variable $X$ denote the number of failures, $X\sim\mathcal{B}(n,1\text{-}\theta)$, where $\theta=0.5$. An unbiased estimator for $\theta$ is ${\hat{\theta}(X)=(n-X)/n}$. The corresponding estimate is referred to as accuracy. The probability of, say, $18$ correct predictions or more corresponds to $x=n-18$ failures or less, that is, the cumulative distribution $P(X\leq x|n,1\text{-}\theta)$. This corresponds to the cumulative distribution for the estimator $\hat{\theta}(X)=\frac{n\text{-}X}{n}$, and we have that $P(\hat{\theta}(X)\geq\frac{n\text{-}x}{n})=P(X\leq x)=0.00020$ for $\hat{\theta}\geq 0.90$. In other words, if $\theta=0.5$, there is a probability of $0.00020$ that the accuracy is $0.90$ or above in a series of $n=20$ trials. We can define the experiment described above as a Bernoulli trial, where success is defined as $X\leq x$, with corresponding probability $P(X\leq x)$, which we from now on write as $P_{x}$ for ease of notation.
If we have $m$ classifiers, we have a series of $m$ Bernoulli trials, and the number of successes follow the binomial distribution $\mathcal{B}(m,P_{x})$. Let $C_{x}$ be the random variable that denotes the number of successful experiments, that is, the number of classifiers that correctly predicts at least $n-x$ coin flips. Of special interest is the case where at least one experiment succeeds, $P(C_{x}>0)=1-P(C_{x}=0)$.
@@
The true SOTA for coin-flip classifiers is $\theta_{SOTA}=0.5$, since it is impossible to train a classifier to predict coin flips. If $m=1,000$ classifiers independently give their predictions for $n=20$ flips, the probability of at least one classifier achieving an accuracy of at least $0.90$ is $0.181-P(C_{x}=0|m,P_{x})=$ 0.18. This means that using the top-ranked method as an estimate and ignoring multiplicity there is a substantial risk of grossly overestimating the SOTA.
The same mechanism of multiplicity applies when $m$ teams submit their classifier’s result on a test set, although in a more complex manner. We seek to describe the probability distribution for increasing levels of complexity, and by that offer a tool to avoid overestimation. We will in the following use the public competition as a general example, and refer to the teams’ classifiers simply as classifiers, for ease of notation.
3.3 The probability distribution of $\hat{\theta}^{\,\max}_{SOTA}(\mathbf{X})$
We define the SOTA probability of success, $\theta_{SOTA}$, as the maximum among the $m$ classifiers being evaluated, $\theta_{SOTA}=\max(\Theta)$, $\Theta=\{\theta_{1},\theta_{2},\ldots,\theta_{m}\}$. Oftentimes, we only have access to sample estimates, $\hat{\Theta}=\{\hat{\theta}_{1},\ldots,\hat{\theta}_{m}\}$. Let $\operatorname{\hat{\theta}^{\,max}_{SOTA}}(\mathbf{X})=\max(\{\hat{\theta}_{1}(X),\hat{\theta}_{2}(X),\ldots,\hat{\theta}_{m}(X)\})$ be the naive, sample maximum estimator for $\theta_{SOTA}$. In the coin flip classifier example above, we have that $\theta_{SOTA}=0.5$, but $P(\hat{\theta}^{\,\max}_{SOTA}(\mathbf{X})\geq 0.90)=0.18$, and it is clear that we risk an overoptimistic estimate.
Of interest is $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}}(\mathbf{X})|m,n,\Theta)$, which we will denote by $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ for ease of notation. Similarly, we use $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ for $\operatorname{\hat{\theta}^{\,max}_{SOTA}}(\mathbf{X})$ when there is no risk of confusion. We here give $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ in the simple case of independent classifiers with identical probabilities of success: $\theta_{j}=\theta$ for all classifiers, $j=1,\ldots,m$. We discuss non-identical probabilities and dependency in Section 4.